OpenStudy (anonymous):

If the ratio in the class of boys to girls is 5 to 10, then how many boys are there in a class of 30 kids?

OpenStudy (curry):

howcome

OpenStudy (anonymous):

or am i wrong??

OpenStudy (anonymous):

let girls be 5x boys = 10x

OpenStudy (anonymous):

5 boys ten girls, or five boys 15 kids. so 30 kids 10 boys

OpenStudy (anonymous):

total = 30 so 5x+10x = 30

OpenStudy (anonymous):

for 15 students there are 5 boys and 10 girls for 30 students there are 10 boys and 20 girls

OpenStudy (curry):

imran got it right

OpenStudy (anonymous):

sry i read wrong

OpenStudy (anonymous):

i read girl to boy is 5 10

OpenStudy (anonymous):

6 boys

OpenStudy (amistre64):

the ratio is 1:2 :) 1:2 = 3 5(1:2 = 3), 5:10 = 15 10(1:2 = 3), 10:20 = 30

OpenStudy (anonymous):

what????

OpenStudy (anonymous):

ya amistre is right

OpenStudy (anonymous):

10 are viewing this

OpenStudy (anonymous):

yup

OpenStudy (curry):

so wats the right answer

OpenStudy (curry):

so 20 right?

OpenStudy (amistre64):

lol ... the right answer is in the journey ;)

OpenStudy (anonymous):

lol

OpenStudy (across):

imran is correct. In fancy formulæ: We know that for every 10 girls, there are 5 boys: \[10b=5g\]\[g=2b\]A classroom composed of girls and boys makes up a total of 30 kids: \[g+b=30\]By substituting: \[2b+b=30\]\[3b=30\]\[b=10\]Therefore, there are 10 boys in the classroom. QED

OpenStudy (shadowfiend):

Bonus points for using the æ ligature, across :p